# Uncovering Semantic Selectivity of Latent Groups in Higher Visual Cortex with Mutual Information-Guided Diffusion

### v1

@@ 3.3 Semantic Image Editing with Deterministic DDIM
Formally, we employ a two-stage deterministic image synthesis process. First, we take an original image $\mathbf{y}_{0}$ and apply deterministic DDIM Inversion (Song et al., 2020a; Mokady et al., 2023) using the diffusion denoiser $\bm{\epsilon}_{\theta}(\cdot)$ from $t=0$ up to a predefined timestep $t=t^{\prime}$. This inversion timestep is calibrated to corrupt the semantic attributes of the original image while preserving its structural information. In the second stage, we reverse the process from $t=t^{\prime}$ to $t=0$ using a classifier-guided, deterministic DDIM sampling. This reverse procedure is guided by our proposed mutual information (MI) maximization objective, ultimately yielding the synthesized image. This overall process is illustrated in Fig. 2(A). The complete synthesis procedure and a detailed algorithm are provided in Appendix B.
@@ 4 Experiments
4.1 Setup
Macaque IT Cortex Dataset. We use single-unit spiking responses from the IT cortex of two macaques (denoted as M1 and M2) during a passive object recognition task (Majaj et al., 2015). Each head-fixed macaque passively viewed a stream of grayscale naturalistic object images while electrophysiological activity in the IT cortex was recorded. Neural activity was recorded from 110 IT channels in M1 and 58 channels in M2. Each image was presented for 100 ms at the center of gaze, and neural responses were measured in a post-stimulus window of 70–170 ms. The stimulus set consisted of 5760 images spanning eight basic-level categories, as illustrated in Fig. 3. To introduce identity-preserving variation, images were simulated via ray-tracing with randomized object position, scale, and pose. As previous studies (Lindsey & Issa, 2024) have shown that IT neurons exhibit a degree of invariance to background context, we segment out the foreground object in each stimulus image.
@@ 4.1 Setup
Figure 4: Synthesized images by MIG-Vis under varying guidance strengths. (A) Results using a frontal-view face and table as the original image, guiding toward a pear. (B) Results using a car and hedgehog object as the original image, guiding toward a plane.
@@ 4.2 Diffusion Visualization of Semantic Selectivity in Latent Groups
To conduct the diffusion visualization with MIG-Vis, we select two key components: a reference image $\mathbf{y}_{0}$ that serves as the base for the editing process described in Section 3.3, and a specific neural latent group $\mathbf{z}_{g}$ from a target image, which serves as the target latent group for our MI-guided synthesis. We visualize images synthesized by MIG-Vis under varying guidance strengths, $\gamma\in\{2,5,10\}$. In Fig. 4(A), a frontal-view face and a table are used as the reference images. Both of them are guided towards the neural latent groups of a target rotated pear. In Fig. 4(B), a car and a hedgehog serve as the reference images, and both are guided toward the neural latent groups of a target plane. Our visual-semantic selectivity findings are as follows:
(1) Latent Group 1: Intra-Category Pose. Probing Latent Group 1 modulates intra-category, pose-related features, for instance the rotation of the face and car objects. This finding is consistent with the rotation supervision applied to this group’s dimensions. Crucially, the objects’ category remain stable throughout all these synthesis, indicating that this latent group disentangles pose variations from semantic content.
(2) Latent Group 2: Inter-Category Semantic Attributes. Notably, despite being supervised only with high-level category identity and no explicit semantic features, Latent Group 2 learns to control inter-category semantic attributes. For example, our MI-guidance transforms the face image into a strawberry and subsequently into a pear. We also observe a direct relationship between activation strength and semantic distance; a larger magnitude of $\gamma$ results in a generated image that is more semantically distinct from the original, demonstrating that the axes of this group subspace encode high-level categorical information.
(3) Latent Group 3 and 4: Intra-Category Content Details. Latent Groups 3 and 4, which were discovered without supervision, both encode intra-category content variations. Specifically, Group 3 primarily modulates the appearance of the face and the chair, whereas Group 4 significantly alters the car and hedgehog’s appearance. Importantly, we note that the selectivity of these two groups to intra-category content details is discovered without supervision. Moreover, Group 3 primarily modulates the appearance of the face with minimal effect on the car. Conversely, Group 4 significantly alters the appearance of the car and hedgehog while inducing little change in the face. This emergent separation likely reflects the distinct visual patterns of variation inherent to these two categories.
Furthermore, as shown in both Fig. 4 (A) and (B), when different original images are guided by the latent group of the same reference image, the synthesized results exhibit consistent semantic variations. This observation verifies the efficacy of our MI guidance objective. Additional visualizations using other reference objects or categories are provided in the Appendix C.
@@ 4.3 Neural Guidance Baseline Methods for Comparison
[…]
Comparison Results Analyses. In Fig. 5, we compare MIG-Vis with the four baselines introduced above, using a frontal-view face as the reference image. We focus on two representative latent groups for this analysis: Group 1 and Group 2, which have been identified as controlling pose and inter-category factors, respectively. We first find that images manipulated by SLT fall outside the natural image distribution and are structurally blurry. AP-LR suffers from a similar limitation, as its gradient is derived from a limited linear model. AP-CFG performs reasonably well in capturing inter-category variations, Nonetheless, it is incapable of uncovering rotation variations due to the entangled gradient of the denoiser and the guidance. Importantly, for the ablation of our framework (w/o MI guidance), we find that it captures rotation semantics effectively. However, for inter-category semantic features, due to the complex underlying structure of this latent subspace, linear activation probing alone is insufficient. This leads to synthesized results that are inconsistent with the realistic categorical variations. In contrast, our MI guidance objective is capable of learning inter-category variations and conducting smooth transitions across object instances.
Figure 6: Pair-wise visual-semantic feature investigation within latent Group 3. (A) Synthesized images of MIG-Vis from probing a frontal-view human face. (B) Synthesized images of MIG-Vis from probing a table object. These results demonstrate that various dimension pairs within a latent group further encode distinct and fine-grained visual-semantic attributes.
@@ 4.4 Fine-Grained Neural Selectivity within Latent Group
In Figures 6 and 7(A), we further investigate the semantic structure of the group subspace by probing single-dimension or pair-wise dimensions within the group. Focusing on Group 2 (“inter-category”) and Group 3 (“intra-category”), we find that these pair-wise subspaces also encode distinct semantic attributes. For example, within the face subspace of Group 3, shown in Fig. 6(A), different dimension pairs control specific visual features: pair $[1,4]$ alters the gaze direction, pair $[3,5]$ modulates the formation of facial features.
An interesting finding from our latent dimension manipulations is that certain combinations exhibit object-invariant effects, while others are object-specific. For instance, the two-dimensional subspace defined by dimensions $[3,5]$ appears to encode a consistent visual-semantic variation across both face and table objects, one that progressively removes fine-grained detail to produce a more abstract, simplified form. In contrast, other combinations of dimensions, such as those involving dimensions $[2,4,6]$, encode distinct, object-specific semantics. For the face, this subspace primarily governs pose and rotation. For the table, however, the same subspace modulates surface properties, including color and the thickness of its components.
We think that semantic probing with MIG-Vis can provide insights into the high-dimensional neural space of the higher visual cortex. One hypothesis is that different latent subspaces encode objects and semantic attributes differently. For example, in Fig. 7(B), on a neural manifold defined by dimensions $[3,5]$, the encodings of “table” and “face” may be close, so MI-guided traversal to “pear” follows similar paths, yielding similar semantic transitions. In contrast, on the manifold of dimensions $[2,4,6]$ shown in Fig. 7(C), “table” and “face” are far apart, producing divergent paths to “pear” and thus distinct semantic variations. This represents just one possible hypothesis. With further MIG-Vis manipulations across more objects and dimensions, we can build a richer geometric intuition of how objects are positioned in neural subspaces, their relative distances, and the semantic attributes each area on the neural manifold encodes. MIG-Vis serves as an intuitive tool for visualizing neural manifolds and generating hypotheses, while also pointing toward future neuroscience research on formally characterizing the geometry of neural subspaces in higher visual cortex.
@@ 4.5 Neural Reconstruction Evaluation
To verify that our neural VAE module maintains high-quality neural reconstruction with the partial correlation regularization and weak label supervision, we quantitatively evaluated its performance. Table 1 presents the R-squared ($R^{2}$ in %) values for both macaques, comparing our module to standard unsupervised VAE. These results demonstrate that the neural reconstruction quality is well-preserved, incurring only a marginal performance drop compared to the standard VAE. We hypothesize that the weakly-supervised labels induce a rotation of the neural latent subspace while preserving the information necessary for reconstruction. From the $R^{2}$ results of ablation studies on our VAE module’s two introduced terms, we observe that each has only a minimal negative impact on neural reconstruction.
Figure 7: (A) Pair-wise visual-semantic feature investigation within latent Group 2. Results of MIG-Vis from probing a table object. (B) Neural Manifolds of Pair-wise Dimensions.
@@ 5 Conclusion
Appendix A Details of the MIG-Vis Implementation
We provide the code and dataset used in this work at the following repository: https://github.com/yulewang97/MIG-Vis. The same codebase is also included in the supplementary ZIP file for convenience.
@@ A.1 Model Training Details
For the partially disentangled neural VAE, we employ an architecture consisting of a two-layer MLP for both the probabilistic encoder and the probabilistic decoder networks. Each MLP uses ReLU activations. For the neural latent dimension number $k$ to both the two macaques, we set $k=24$ and partition it into four distinct latent groups, each of dimension $6$ (i.e., group rank equals to $6$). This group rank setting provides the expressiveness to model complex visual semantic factors (e.g., inter-category variations) within a unique latent group. For latent group $1$, we weakly supervise its learning with the 3D rotation angles of the object provided in the dataset; for latent group 2, we use the 8-way category identity labels (e.g., “Animals” as integer label 0, “Boats” as integer label $1$); latent groups 3 and 4 are learned in a fully unsupervised manner. We use a learning rate of $1\times 10^{-3}$ and set partial correlation penalty pc_weight to $5\times 10^{-5}$ on both two macaques. We use the Adam Optimizer Kingma & Ba (2014) for optimization.
@@ Appendix B Detailed Mutual-Information Maximization Guidance Algorithm
Input: Target Neural Latent Group $\mathbf{z}_{g}$, Guidance Scale $\gamma$, Sampling Step $t^{\prime}$
@@
8 $\hat{\bm{\epsilon}}(\mathbf{y}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)+\gamma\,\nabla_{\mathbf{y}_{t}}\mathcal{L}_{\mathrm{N}}\left(\bm{\phi}\right)$ $\triangleright$ Mutual Information Guided Gradient ;
@@
Appendix C Additional Visualizations of Disentangled Semantic Latent Groups
Figure 8: Synthesized images by MIG-Vis under varying guidance strengths. (A) Results using a watermelon and a peach as the reference image, guiding toward a rotated face. (B) Results using a ship and a airplane object as the reference image, guiding toward a strawberry. Group 1 captures pose-related variations (e.g., object rotation), Group 2 controls inter-category semantic transformations, while Groups 3 and 4 govern distinct intra-category content variations. Image transitions along guidance strength $\gamma$ demonstrate smooth and interpretable semantic information.
Figure 9: Synthesized images by MIG-Vis under varying guidance strengths. (A) Results using a face and a berry as the reference image, guiding toward a rotated table. (B) Results using a chair and a boat object as the reference image, guiding toward a car. Group 1 captures pose-related variations (e.g., object rotation), Group 2 controls inter-category semantic transformations, while Groups 3 and 4 govern distinct intra-category content variations. Image transitions along guidance strength $\gamma$ demonstrate smooth and interpretable semantic information.

### v2

@@ 3.3 Semantic Image Editing with Deterministic DDIM
Formally, we employ a two-stage deterministic image synthesis process. First, we take an original image $\mathbf{y}_{0}$ and apply deterministic DDIM Inversion (Song et al., 2020; Mokady et al., 2023) using the diffusion denoiser $\bm{\epsilon}_{\theta}(\cdot)$ from $t=0$ up to a predefined timestep $t=t^{\prime}$. This inversion timestep is calibrated to corrupt the semantic attributes of the original image while preserving its structural information. In the second stage, we reverse the process from $t=t^{\prime}$ to $t=0$ using a classifier-guided, deterministic DDIM sampling. This reverse procedure is guided by our proposed MI maximization objective, ultimately yielding the synthesized image. This overall process is illustrated in Fig. 2(A). The complete synthesis procedure and a detailed algorithm are provided in Appendix B.
@@ 4 Experiments
Macaque IT Cortex Dataset. We use single-unit spiking responses from the IT cortex of two macaques (M1 and M2) during a passive object recognition task (Majaj et al., 2015). Each head-fixed macaque passively viewed grayscale naturalistic object images while electrophysiological activity was recorded from 58 IT channels in M1 and 110 in M2. Each image was shown for 100 ms at the center of gaze, and neural responses were measured within a 70–170 ms post-stimulus window. The dataset contains 5760 images spanning eight basic-level categories (Fig. 3). To introduce identity-preserving variation, images were generated via ray-tracing with randomized object position, scale, and pose. Following prior work (Lindsey and Issa, 2024), we segment the foreground in each image to reduce background effects.
@@
Figure 4: Synthesized images by MIG-Vis under varying guidance strengths. Results using a frontal-view face, car, strawberry, and table as the original images.
@@ 4.1 Diffusion Visualization of Semantic Selectivity in Latent Groups
To perform diffusion-based visualization with MIG-Vis, we select two key components: an original image $\mathbf{y}_{0}$, which serves as the starting point for the editing process described in Section 3.3, and a neural latent group $\mathbf{z}_{g}$, which we perturb for MI-guided synthesis. We generate images using MIG-Vis under different guidance strengths, $\gamma\in\{-10,-5,5,10\}$. In Fig. 4, a frontal-view face, a car, a strawberry, and a table are used as the original images. For each case, we perturb the corresponding neural latent with both positive and negative $\gamma$. The resulting visual-semantic selectivity is summarized as follows:
(1) Latent Group 1: Pose. Probing Latent Group 1 primarily modulates pose-related features, such as the rotation of faces and cars. This aligns with the rotation supervision applied to this group’s dimensions. Importantly, object category remains unchanged across all synthesized images, indicating that this latent group separates pose variation from other semantic content.
(2) Latent Group 2: Inter-Category Semantic Attributes. Notably, despite being supervised only with high-level category identity and no explicit semantic features, Latent Group 2 learns to control inter-category semantic attributes. For example, our MI-guidance transforms the face image into a strawberry. We also observe a direct relationship between activation strength and semantic distance; a larger magnitude of $\gamma$ results in a generated image that is more semantically distinct from the original, demonstrating that the axes of this group subspace encode high-level categorical information.
(3) Latent Group 3 and 4: Intra-Category Content Details. Latent Groups 3 and 4, discovered without supervision, both encode intra-category content variations. Group 3 primarily modulates the appearance of faces and strawberries with little effect on cars, whereas Group 4 significantly alters cars and tables while minimally affecting faces. This selective behavior likely reflects distinct patterns of visual variation across object categories. These results suggest that the neural latent manifold in these dimensions is locally structured rather than globally aligned. Different object categories occupy different regions of the manifold and vary along different directions. Geometrically, this implies that the manifold is anisotropic and curved: the principal directions of variation are category-dependent. Faces and strawberries extend strongly along the dimensions of Group 3, whereas cars and tables extend more along Group 4. Thus, intra-category variation is organized into separate local tangent directions rather than a single shared global axis across all objects.
@@ 4.2 Neural Guidance Baseline Methods for Comparison
Comparison Results Analyses. In Fig. 5, we compare MIG-Vis with the three baselines introduced above, using a frontal-view face as the original image. We focus on two representative latent groups for this analysis: Group 1 and Group 2, which have been identified as controlling pose and inter-category factors, respectively. We first observe that images manipulated by SLT exhibit some rotation when probing Group 1, but the changes are not as clean as those produced by other methods. When probing Group 2, the object category remains unchanged. This suggests a limitation of decoder-based image generation compared to diffusion-based generation. AP-CFG performs reasonably well in capturing rotation semantics; however, its perturbations for inter-category variations are less clean compared to our MIG-Vis.
Importantly, in the ablation of our framework w/o MI guidance, where diffusion is guided by likelihood-based alignment $\nabla_{\mathbf{y}}\log p(\mathbf{z}_{g}|\mathbf{y})$, we observe that pose-related rotation semantics can be captured moderately well. Both ours with and without MI rely on linear activation probing to define the latent subspace. However, the key difference lies in the guidance objective. Likelihood-based guidance only enforces that the synthesized image be mapped back to the target latent $\mathbf{z}_{g}$ by the learned probabilistic encoder. In other words, diffusion merely needs to generate an image that the encoder recognizes as having latent $\mathbf{z}_{g}$. This constraint is one-sided and encoder-dependent: as long as the encoder assigns high likelihood to $\mathbf{z}_{g}$, the objective is satisfied. For simple, low-dimensional variations such as rotation, this may be sufficient. However, for inter-category semantic features, whose latent structure is complex and nonlinear, this objective becomes too weak. Multiple visually different images may map to similar latent values under the encoder, and subtle semantic structure may be averaged out. As a result, likelihood-based guidance can produce inconsistent or unrealistic categorical transitions.
In contrast, our MIG-Vis employs MI guidance, which imposes a stronger constraint. Rather than asking whether the encoder can recognize the image as having latent $\mathbf{z}_{g}$, MI maximization requires that the synthesized image and $\mathbf{z}_{g}$ share maximal statistical dependence. Intuitively, likelihood guidance asks: Can the encoder recognize this image as having latent $\mathbf{z}_{g}$? MI guidance asks: Does this image truly express the information contained in $\mathbf{z}_{g}$? The latter is a stricter requirement. It forces the diffusion process to generate images that visibly and structurally reflect the semantic content encoded in the perturbed latent group, rather than merely satisfying an encoder consistency check. Consequently, MI guidance better preserves complex inter-category structure and produces smooth, realistic transitions across object instances.
Figure 6: Pair-wise visual-semantic probing in latent Group 3. (A) MIG-Vis synthesis from a frontal-view face. (B) MIG-Vis synthesis from a strawberry. These results show that different dimension pairs within the same latent group encode distinct, fine-grained visual-semantic attributes.
@@ 4.3 Fine-Grained Neural Selectivity within Latent Group
Figure 7: (A) Pair-wise visual-semantic feature investigation within latent Group 2. Results of MIG-Vis from probing a table object. (B) and (C): Neural manifolds of two latent groups.
In Figs. 6 and 7(A), we probe pair-wise latent dimensions within the Group 3 intra-category subspace by starting from both a face image and a strawberry image, and within the Group 2 inter-category subspace by starting from a table image. We observe that perturbing the same dimension pairs results in different semantic changes across object categories.
In Fig. 6, for dimension pair $[1,2]$, perturbations applied to the face primarily modulate orientation and gaze direction (e.g., upward gaze or left orientation shift). In contrast, perturbing the same dimensions for the strawberry mainly produces texture smoothing and global lighting changes. Similarly, within the $[3,5]$ subspace, faces exhibit structural variations such as right-side gaze shifts, whereas strawberries show intra-category changes and surface darkening. For $[4,6]$, perturbations deform facial features and alter viewpoint in faces, but induce shape flattening and stripe variations in strawberries.
With these semantic probing visualizations produced by MIG-Vis, we gain insight into the structure of the high-dimensional neural space in the higher visual cortex. Here, we illustrate plausible neural manifolds within example subspaces.
For the latent dimensions associated with Group-1 (pose), increasing $\gamma$ produces systematic rotational changes across all object categories (Fig. 4). However, the direction of rotation depends on the object: faces rotate clockwise while cars rotate counterclockwise. A similar pattern is observed for strawberries and tables. This indicates that perturbations along these latent dimensions consistently encode a rotation semantic, regardless of object identity. In other words, manipulating the same latent axis always induces rotation, even though the visual manifestation of that rotation may differ across categories. This observation suggests that the underlying geometry of the latent space in these dimensions is structured rather than linear. A natural hypothesis is that the manifold along these dimensions resembles a torus (Fig. 7B). In this view, different object categories occupy different locations on the toroidal manifold: cars and tables lie closer on one side, while faces and strawberries lie closer on the opposite side. When a latent perturbation is applied (vertical arrows in Fig. 7B), its projection onto the toroidal manifold induces motion along the surface in opposite directions for different object regions. This results in clockwise rotation for some objects and counterclockwise rotation for others. Crucially, despite these differences in local behavior, the semantic meaning of movement along this latent axis remains consistent across categories: it corresponds to object rotation. We therefore interpret these dimensions as forming a globally consistent semantic manifold in neural latent space.
In contrast, the latent dimensions associated with Group-3 (intra-category) exhibit a fundamentally different behavior. For example, when perturbing a face along dimension 1 versus dimension 2 with changing $\gamma$, we observe distinct semantic changes such as gaze direction and feature formation. However, applying the same perturbations to a different object, such as a strawberry, produces entirely different semantic variations, such as changes in texture smoothing or stripe patterns. This stands in sharp contrast to the Group-1 pose dimensions, where latent perturbations consistently correspond to the same semantic transformation across objects. Here, the same latent directions do not induce a shared semantic effect when applied to different object categories. This suggests that the underlying manifold in this subspace is substantially more complex and nonlinear (e.g., Fig. 7C). Rather than forming a globally structured surface like the toroidal rotation manifold in Group-1, the geometry here appears warped and irregular. Perturbations along different latent axes project onto the manifold in different ways: some may induce motion along a locally structured trajectory (e.g., along an “8-shaped” curve), while others act in directions that are effectively orthogonal to it. As a result, the semantic interpretation of these dimensions becomes object-dependent. Unlike the globally consistent rotation semantics observed in Group-1, no shared transformation emerges across categories. Instead, semantic meaning can only be interpreted locally, relative to each object’s position on the manifold. Taken together, this indicates that while some latent subspaces encode globally consistent semantics (e.g., pose), others correspond to highly warped intra-category manifolds whose effects are inherently local rather than universal.
Thus, MIG-Vis serves as an intuitive tool for visualizing neural manifolds and generating hypotheses, while also pointing toward future neuroscience research on formally characterizing the geometry of neural subspaces in higher visual cortex.
@@ 4.4 Neural Reconstruction Evaluation
To verify that our neural VAE module maintains high-quality neural reconstruction with the partial correlation regularization and weak label supervision, we quantitatively evaluated its performance. Table 1 presents the R-squared ($R^{2}$ in %) values for both macaques, comparing our module to standard unsupervised VAE. These results demonstrate that the neural reconstruction quality is well-preserved, incurring only a marginal performance drop compared to the standard VAE. We hypothesize that the weakly-supervised labels induce a rotation of the neural latent subspace while preserving the information necessary for reconstruction. From the $R^{2}$ results of ablation studies on our VAE module’s two introduced terms, we observe that each has only a minimal negative impact on neural reconstruction. We also plot the raw neural signals and their reconstruction across dimensions for four selected trials in Fig. 8 in the Appendix A.4.
@@ 5 Conclusion
Acknowledgement
This work was supported by the seed grant from Microsoft GenAI.
@@ Acknowledgement
Appendix A Details of the MIG-Vis Method
@@ A.1 Model Training Details
For the partially disentangled neural VAE, we employ an architecture consisting of a two-layer MLP for both the probabilistic encoder and the probabilistic decoder networks. Each MLP uses ReLU activations. For the neural latent dimension number $k$ to both the two macaques, we set $k=24$ and partition it into four distinct latent groups, each of dimension $6$ (i.e., group rank equals to $6$). This group rank setting provides the expressiveness to model complex visual semantic factors (e.g., inter-category variations) within a unique latent group. For latent group $1$, we weakly supervise its learning with the 3D rotation angles of the object provided in the dataset; for latent group 2, we use the 8-way category identity labels (e.g., “Animals” as integer label 0, “Boats” as integer label $1$); latent groups 3 and 4 are learned in a fully unsupervised manner. We use a learning rate of $1\times 10^{-3}$ and set partial correlation penalty pc_weight to $5\times 10^{-5}$ on both two macaques. We use the Adam Optimizer Kingma and Ba (2014) for optimization.
@@ A.3 Group-wise Disentangled Neural Latent Subspace Investigation
Table 3: Explained variance of neural signals from the supervised and unsupervised neural latents of the two macaques M1 and M2.
Subject Metrics Full Latents Supervised Latents Unsupervised Latents
M1 $R^{2}$ (%) $\uparrow$ $76.13\,(\pm 0.27)$ $32.50\,(\pm 0.19)$ $20.88\,(\pm 0.25)$
RMSE $\downarrow$ $48.86\,(\pm 0.21)$ $0.75\,(\pm 0.28) 88$.95(±0.29)88.95\,($\pm$ 0.29)
M2 $R^{2}$ (%) $\uparrow$ $71.82\,(\pm 0.28)$ $28.11\,(\pm 0.22)$ $19.63\,(\pm 0.17)$
RMSE $\downarrow$ $53.09\,(\pm 0.16)$ $84.79\,(\pm 0.32)$ $89.65\,(\pm 0.31)$
A.4 Neural signal reconstruction
Figure 8: (A) Raw Neural signals and the reconstructions from the neural LVM module in MIG-Vis on the M1 subject. (B) Raw Neural signals and the reconstructions from the neural LVM module in MIG-Vis on the M2 subject.
Figure 9: Neural latent group number $G$ analyses. (A) Synthesized images of MIG-Vis from probing a frontal-view human face. (B) Synthesized images of MIG-Vis from probing a car object.
Figure 10: InfoNCE negative sample $B$ analyses in Group-1 (Pose). (A) Synthesized images of MIG-Vis from probing a frontal-view human face. (B) Synthesized images of MIG-Vis from probing a car object.
@@ Appendix B Detailed Mutual-Information Maximization Guidance Algorithm
Input: Target Neural Latent Group $\mathbf{z}_{g}$, Guidance Scale $\eta$, Sampling Step $t^{\prime}$
@@
8 $\hat{\bm{\epsilon}}(\mathbf{y}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)+\eta\,\nabla_{\mathbf{y}_{t}}\mathcal{L}_{\mathrm{N}}\left(\bm{\phi}\right)$ $\triangleright$ Mutual Information Guided Gradient ;